# Solving Dicke superradiance analytically: A compendium of methods

### v1

@@ Abstract
We revisit the Dicke superradiance problem, namely the time evolution of the density operator for an initially inverted ensemble of $N$ identical two-level systems (spins), undergoing collective spontaneous emission. This is one of the simplest examples of an open quantum system dynamics problem that admits a fully analytical solution. We provide a few different approaches to derive the analytical solution valid for any time during the evolution and for any number of spins. These approaches range from solving a set of coupled rate equations, to observing exceptional points stemming from non-Hermitian evolution, to the use of combinatorics and probabilistic methods, and finally to utilizing techniques based on the quantum jump unraveling of the master equation. The solution is expressed as a residue sum stemming from a contour integral in the complex plane and suggests the possibility of fully analytical solutions for a wider range of quantum open system dynamics problems.
Dicke superradiance refers to the rapid burst of radiation from an initially fully inverted ensemble of $N$ collectively decaying two-level quantum systems (TLS). Let us denote the spontaneous emission rate of an isolated TLS with $\Gamma$, describing the incoherent loss of excitation from initial excited state |$\ket{e}$ to the ground state |$\ket{g}$ (via a collapse operator $\sigma=\ket{g}\bra{e}$ |). It follows that the fully inverted system |$\ket{ee...e}$, where the TLS are assumed to be at the same position, relaxes back to the collective ground state |$\ket{gg...g}$ with a time-dependent rate. For the initial state, the rate is $\Gamma$, as the emitters start emitting independently. The total radiated power is $N\Gamma$ as there are $N$ emitters. As the system evolves to lower occupancy states, the emission rate increases to a maximum of $N\Gamma$, while the excitation decreases from $N$ to zero. The peak of radiated power appears when the system passes through half-inverted states, where the population is around $N/2$ and the rate achieves close to the maximum rate of $N\Gamma$: the peak of radiated power than shows the superradiance signature of scaling $\propto N^{2}$. The key point of the problem is the presence of a single decaying channel via a collective operator $S=\sum_{j=1}^{N}\sigma_{j}$, which renders the particles indistinguishable. This feature was introduced by Dicke in 1954 [1] and was first analytically solved by Lee in 1977, starting with the full excitation case [2] and generalizing to any arbitrary initial state [3]. Other analytical treatments followed, for example described in Refs. [4, 5].
@@
The solution proposed in Refs. [2, 3] has been mostly overlooked during the last decades; presumably the convoluted analytical expressions renders this solution computationally demanding, thus requiring resources similar to the direct simulation of the master equation, which is a trivial task for particles numbers up to around $N=10^{4}$. The focus instead shifted more recently to derive more informative solutions in the mesoscopic (or thermodynamic) limit of large $N$, such as, for example, in Refs. [6, 7, 8, 9, 10, 11]. A comprehensive early review of the problem and its implications can be found in Ref. [12].
@@
The original formulation of the Dicke superradiance model is hard to realize in practice, as closely positioned quantum emitters interact strongly (proportionally to $d^{-3}$, where $d$ is their separation), shifting the bare electronic levels and removing their indistinguishability. However, alternative realizations are available, for example in the context of cavity quantum electrodynamics [13, 14, 15, 16]. Here, atoms can be trapped in the standing wave field of a lossy optical cavity (with atom-field coupling rate $g$ and photon loss rate $\kappa$) at equivalent positions, i.e., with equal couplings to the light mode. The lossy cavity opens an additional decay channel along the cavity direction. This occurs in the limit of $\kappa\gg g$, where the excitation of the atoms is taken by the cavity field and emitted quickly through the side mirrors. This new decay channel is added to the intrinsic spontaneous emission into all three directions of the free space and is characterized a collective collapse operator $S$ (after adiabatic elimination of the cavity mode) and a collapse rate $\Gamma=g^{2}/\kappa$. Assuming that the cavity induced decay is much larger than the individual atom spontaneous emission, the dynamics of the atoms are fully described the master equation in Eq. (1). This is referred to as cavity-induced superradiance and has been implemented to realize a form of superradiant lasing [17, 18, 19]. Moreover, waveguide and circuit quantum electrodynamics provide alternative, easier to manipulate, platforms to realize Dicke superradiance. In these cases as it is easier to have design ensembles of practically non-interacting artificial electronic systems, placed within a wavelength of the resonator mode [20, 21].
@@
To be more precise, at the mathematical level, the Dicke superradiance problem requires one to find the time evolution of density operator $\rho$ of the $N$ two-level systems undergoing collective decay. The density operator satisfies the following master equation, written in Lindblad form (and here simplified as it is cast in an interaction picture, where the trivial action of the free Hamiltonian is removed):
@@
A system of $N$ two-level system spans a Hilbert space of dimensions of $2^{N}$; the symmetric form of the collapse operator $S$, however, restricts the evolution solely within the symmetric subspace spanned by $N+1$ collective Dicke states |$\ket{m}$, as the initially inverted state is part of this subspace. The index runs over $m=0,1,\ldots,N$, with the fully ground |$\ket{0}=\ket{gg...g}$ and fully excited state |$\ket{N}=\ket{ee...e}$, easily expressed in terms of the bare basis. The symmetric operators $S$ and $S^{\dagger}$ have the following action on the Dicke states: $S|m\rangle=\sqrt{h_{m}}|m-1\rangle$ and $S^{\dagger}|m\rangle=\sqrt{h_{m+1}}|m+1\rangle$, keeping the evolution completely inside the symmetric subspace. The coefficients are defined as $h_{m}=m\bar{m}$ with $\bar{m}=(N+1)-m$ and exhibit a double degeneracy for even $N$ as $h_{m}=h_{\bar{m}}$. For odd $N$, there is an additional non-degenerate point at $(N+1)/2$. The separation of the Lindblad term into two superoperators is based on the observation that $B[\rho]=\Gamma S\rho S^{\dagger}$ lowers the excitation number by one, while $C[\rho]$ conserves the number of excitations in the system.
There are a few competing approaches to find the solution to the problem listed above. We present here five of such approaches, and stress their advantages and shortcomings. For the first four approaches, we will recast the problem into its formal solution as an infinite time series expansion
@@
Note that the repeated action of the superoperator $\mathcal{L}$ onto the initial density operator $\rho_{0}=\rho(0)=\ket{N}\bra{N}$ | is to be carefully evaluated. We remark that the density operator at any time is diagonal in the Dicke basis: this can be seen as noticing the action of the two superoperators $B[\ket{m}\bra{m}]=\Gamma h_{m}\ket{m-1}\bra{m-1}$ | and $C[\ket{m}\bra{m}]=-\Gamma h_{m}\ket{m}\bra{m}$ |, which indicates that no off-diagonal elements of the density operator can be generated during this incoherent evolution, when starting in a Dicke state. This allows to write $\rho(t)=\sum_{m=0}^{N}\rho_{m}(t)\ket{m}\bra{m}$ |, meaning that the task is reduced to finding the fully analytical expression of any $\rho_{m}(t)$ for any time.
@@
    $\rho_{m}(t)=\langle m|e^{\mathcal{L}t}\rho_{0}|m\rangle=\sum_{j=0}^{\infty}\frac{(\Gamma t)^{j}}{j!}\rho_{m}^{(j)}.$ (4)
    The recursion connects the $n+1$ component in the time series of state $m-1$ to the $n$ components stemming either ”longitudinally” from the same state $m-1$ or ”diagonally” from the state situated immediately above $m$. We then construct a solution $\rho^{(n)}_{m}$ for any order $n$ and any state $m$ and sum it up from 00 to infinity, to construct the density operator as an infinite time series $\rho_{m}(t)=\sum_{n=0}^{\infty}t^{n}/n!\rho^{(n)}_{m}$.
@@
    Spelling out the exact contributions of each term in the expansion, leads to a sum over all possible paths, which we then add to find the exact expression for any $\rho_{m}(t)$.
@@
    Probabilistic approach: We expand $e^{\mathcal{L}\Delta t}$ for small $\Delta t$ in order to connect the density operator between discretized moments $j$ and $j+1$, with $t=j\Delta t$ to $t+\Delta t=(j+1)\Delta t$. Two complementary events emerge, one quantified by $d_{j}$, which is the probability to perform a jump within the time interval $\Delta t$ and $s_{j}=1-d_{j}$, quantifying the probability to remain in the respective quantum state. Summing up all possible paths leading from the initially inverted state $\rho_{0}$ to a target state $m$, leads to an analytical expression for $\rho_{m}(t)$.
@@
    Non-Hermitian Hamiltonian approach: We start again with the recursive equations and cast it into a form involving a non-Hermitian Hamiltonian $H$. With this, the general solution involves a matrix exponential requiring a Jordan decomposition. Due to the symmetry $h_{m}=h_{\bar{m}}$, $H$ exhibits non-Hermitian degeneracies called exceptional points, which result in a modified decay compared to a non-degenerate system, ultimately leading to Dicke superradiance.
@@ II Quantum trajectories approach
$\rho(t)=\frac{1}{n_{\text{traj}}}\sum_{j}\ket{\psi(t)}_{j}\bra{\psi(t)}_{j},$ (12)
@@
For a given trajectory, the system spends a total time $\tau_{m}$ in state |$\ket{m}$, with $t$ enclosed between $t_{m+1}$ (the time at which the jump from state $m+1$ into the state $m$ occurs) and $t_{m}$ (when the next jump out of the state occurs). An average of all possible trajectories then gives the value of $\rho_{m}(t)$ as the probability that the system finds itself in state $m$ at time $t$. This corresponds to an average of the segment length $\left[\theta(t-t_{m})-\theta(t-t_{m-1})\right]$ (where $\theta$ denotes the Heaviside function) over all possible values of $p_{N}$,…,$p_{1}$. The evolution is visually represented in Fig. 2. The population of the state $\rho_{m}(t)$ can then be computed as
$\rho_{m}(t)=\int_{0}^{1}\text{d}p_{N}\cdots\int_{0}^{1}\text{d}p_{m}\left[\theta(t-t_{m})-\theta(t-t_{m-1})\right],$ (13)
@@
$\theta(x)=\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\frac{\text{e}^{\text{i}s\Gamma x}}{s-\text{i}\epsilon}.$ (15)
We substitute this into Eq. (14) with the replacement $x$ by $t-\sum_{k=m+1}^{N}\tau_{k}$ first, to evaluate $\mathcal{I}_{m+1}$ and then with the replacement $x$ by $t-\sum_{k=m}^{N}\tau_{k}$ to evaluate $\mathcal{I}_{m}$. We distribute the terms with $\tau_{k}$ from the exponent into all integrals over $\tau_{k}$. This allows to compute any of the integrals which lead to the generation of poles of s atat $-\text{i}h_{k}$:
$\int_{0}^{\infty}\text{d}\tau_{k}e^{-\Gamma h_{k}\tau_{k}}e^{-\text{i}s\Gamma\tau_{k}}=-\frac{1}{(\text{i}s+h_{k})\Gamma}.$ (16)
@@
$\mathcal{I}_{m+1}=-\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\left[\prod_{k=m+1}^{N}\frac{(-h_{k})}{(\text{i}s+h_{k})}\right]\frac{\text{e}^{\text{i}t\Gamma s}}{s-\text{i}\epsilon},$ (17)
@@
$\mathcal{I}_{m}=-\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\left[\prod_{k=m}^{N}\frac{(-h_{k})}{(\text{i}s+h_{k})}\right]\frac{\text{e}^{\text{i}t\Gamma s}}{s-\text{i}\epsilon}.$ (18)
@@ III A combinatorics approach
In order to obtain the most general form of the sequences above, we first placed $r$ superoperators $B$ in the sequence (with $r$ varying from 00 to $n$) and added paddings of $C$s in the spaces between the $B$s, with $j_{N}$, $j_{N-1}$,…, $j_{N-r}$ counting the numbers of superoperators $C$ in the sequence. Summing over $r$ from 00 to $n$ and for any possible sequences (or spellings) corresponding to each fixed $r$ gives all possible combinations involved in the binomial expansion.
@@
where $\delta_{mm^{\prime}}$ is the Kronecker delta. Summing up the infinite time series implies the need for performing the constrained sum:
@@
This can be evaluated via the procedure described in App. A, via the identifications $n_{t}=N-m+1$ and $M=n-(N-m)$, using the following generating function
@@
The constrained sum is the coefficient of the $1/z^{n-(N-m)}$ series expansion of $f_{N-m+1}(z)$ and can be computed as the sum of residues of the function $f_{N-m+1}(z)$ multiplied by $e^{z\Gamma t}$, which is exactly the function listed in Eq. (7). Finally, the solution in Eq. (6) is reproduced.

### v3

@@ Abstract
We present several analytical approaches to the Dicke superradiance problem, which involves determining the time evolution of the density operator for an initially inverted ensemble of $N$ identical two-level systems undergoing collective spontaneous emission. This serves as one of the simplest cases of open quantum system dynamics that allows for a fully analytical solution. We explore multiple methods to tackle this problem, yielding a solution valid for any time and any number of spins. These approaches range from solving coupled rate equations and identifying exceptional points in non-Hermitian evolution to employing combinatorial and probabilistic techniques, as well as utilizing a quantum jump unraveling of the master equation. The analytical solution is expressed as a residue sum obtained from a contour integral in the complex plane, suggesting the possibility of fully analytical solutions for a broader class of open quantum system dynamics problems.
Dicke superradiance describes the rapid burst of radiation emitted by an initially fully inverted ensemble of $N$ two-level quantum systems (TLS) undergoing collective decay [1]. This is in stark contrast with the situation exhibited by a dilute, non-interacting, independently emitting ensemble. To understand the emergence of such behavior, let us denote by $\Gamma$ the spontaneous emission rate of an isolated TLS from its excited state |$\ket{e}$ to its ground state |$\ket{g}$. The collapse operator associated with this loss process is denoted by $\sigma=\ket{g}\bra{e}$ |. For independent emitters, the rate at which the ensemble emits is identical to the one of an isolated system, so each constituent of the ensemble will lose excitation exponentially $\propto e^{-\Gamma t}$. Therefore, the time dependence of the total radiated power shows a trivial exponential dependence $\propto N\Gamma e^{-\Gamma t}$. For the collectively decaying system, the collapse operator is $S=\sum_{j=1}^{N}\sigma_{j}$ and the decay rate turns out to be strongly time dependent. The fully inverted state decays at the same rate per particle as the isolated TLS, i.e. at rate $\Gamma$; however, the emergence of particle-particle correlations leads to an increase in the emission rate as the ensembles evolves toward half-occupancy states, to a maximum of $N\Gamma$ per particle. Meanwhile, as the total excitation number decreases from $N$ to zero, a peak of radiated power occurs when the system passes through the half-inverted state and the emission rate is close to the maximum rate of $N\Gamma$ per particle. This peak of total radiated power with amplitude $\propto N^{2}$ and occurring at short times $\propto\ln(N)/(N\Gamma)$ is the typical signature of superradiance [2].
The main aspect of Dicke superradiance is the permutation symmetry, owing to the symmetry of the collective collapse operator $S$. This feature was introduced by Dicke in 1954 [1] and analytically solved first by Lee in 1977, starting with the full excitation case [3] and generalizing to any arbitrary initial Dicke state [4]. Other analytical treatments followed, for example described in Refs. [5, 6].
@@
The solution proposed in Refs. [3, 4] has been mostly overlooked during the last decades. Presumably the convoluted analytical expression listed therein renders this solution computationally demanding, thus requiring resources similar to the direct simulation of the master equation. The focus instead shifted recently to derive more informative solutions in the mesoscopic (or thermodynamic) limit of large $N$, such as, for example, in Refs. [7, 8, 9, 10, 11, 12]. A comprehensive early review of the problem and its implications can be found in Ref. [2].
@@
The original formulation of the Dicke superradiance model is hard to realize in practice, as closely positioned quantum emitters interact strongly (proportionally to $d^{-3}$, where $d$ is their separation), shifting the bare electronic levels and removing their indistinguishability. However, alternative realizations are available, for example in the context of cavity quantum electrodynamics [13, 14, 15, 16]. Here, atoms can be trapped in the standing wave field of a lossy optical cavity (with atom-field coupling rate $g$ and photon loss rate $\kappa$) at equivalent positions, i.e., with equal couplings to the cavity light mode. The lossy cavity opens an additional decay channel along the cavity direction. This occurs in the limit of $\kappa\gg g$, where the excitation of the atoms is taken by the cavity field and emitted quickly through the side mirrors. This new decay channel is added to the intrinsic spontaneous emission into all three directions of free space and is characterized by a collective collapse operator $S$ (after adiabatic elimination of the cavity mode) and an associated collapse rate $\Gamma=g^{2}/\kappa$. Assuming that the cavity induced decay is much larger than the individual atom spontaneous emission, the dynamics of the atoms are fully described the master equation in Eq. (1). This is referred to as cavity-induced superradiance and has been implemented to realize a form of superradiant lasing [17, 18, 19]. Moreover, waveguide and circuit quantum electrodynamics provide alternative, simpler to manipulate, platforms to realize Dicke superradiance. In these platforms, it is easier to design ensembles of practically non-interacting artificial electronic systems, placed at multiples of the wavelength of the resonator mode [20, 21].
@@
Mathematically, the Dicke superradiance problem involves determining the time evolution of the density operator $\rho(t)$ for $N$ two-level systems undergoing collective decay. The density operator follows the master equation in Lindblad form, presented here in a simplified version within the interaction picture, where the trivial effect of the free Hamiltonian has been eliminated:
@@
The separation of the Lindblad term into two superoperators is based on the observation that $B[\rho]=\Gamma S\rho S^{\dagger}$ lowers the excitation number by one, while $C[\rho]$ conserves the number of excitations in the system. The system of $N$ two-level system spans a Hilbert space of dimension $2^{N}$; the symmetric form of the collapse operator $S$, however, restricts the evolution solely within the symmetric subspace spanned by $N+1$ collective Dicke states |$\ket{m}$, as the initially inverted state is part of this subspace. The index runs over $m=0,1,\ldots,N$. The ground |$\ket{0}=\ket{gg...g}$ and fully excited state |$\ket{N}=\ket{ee...e}$, are easily expressed in terms of the bare basis. The symmetric operators $S$ and $S^{\dagger}$ have the following action on the Dicke states
$S|m\rangle$ $=\sqrt{h_{m}}|m-1\rangle$ (2a)
$S^{\dagger}|m\rangle$ $=\sqrt{h_{m+1}}|m+1\rangle,$ (2b)
keeping the evolution completely inside the symmetric subspace. The coefficients are defined as
$h_{m}=m\bar{m}\qquad\text{with}\qquad\bar{m}=(N+1)-m$ (3)
and exhibit a double degeneracy for even $N$ as $h_{m}=h_{\bar{m}}$. For odd $N$, there is an additional non-degenerate point at $m=(N+1)/2$.
There are a few competing approaches to find the solution to the problem listed above. We present here five such approaches, and stress their advantages and shortcomings. For the first four approaches, we will recast the problem into its formal solution as an infinite time series expansion
@@
Thus, it is required to evaluate the repeated action of the superoperator $\mathcal{L}$ onto the initial density operator $\rho_{0}=\rho(0)=\ket{N}\bra{N}$ |. We remark that the density operator at any time is diagonal in the Dicke basis: this can be seen by noticing the action of the two superoperators $B[\ket{m}\bra{m}]=\Gamma h_{m}\ket{m-1}\bra{m-1}$ | and $C[\ket{m}\bra{m}]=-\Gamma h_{m}\ket{m}\bra{m}$ |, which indicates that no off-diagonal elements of the density operator can be generated during this incoherent evolution, when starting in a Dicke state. This allows to write $\rho(t)=\sum_{m=0}^{N}\rho_{m}(t)\ket{m}\bra{m}$ |, meaning that the task is reduced to finding the expression of $\rho_{m}(t)$ for any time.
@@
    $\rho_{m}(t)=\langle m|e^{\mathcal{L}t}\rho_{0}|m\rangle=\sum_{n=0}^{\infty}\frac{(\Gamma t)^{n}}{n!}\rho_{m}^{(n)}.$ (6)
    The recursion connects the $n+1$ component in the time series of state $m-1$ to the $n$ components stemming either ”longitudinally” from the same state $m-1$ or ”diagonally” from the state situated immediately above $m$. We then construct a solution $\rho^{(n)}_{m}$ for any order $n$ and any state $m$ and sum it from zero to infinity, to construct the density operator as an infinite time series.
@@
    Writing out the contributions of each term in the expansion leads to a sum over all possible paths, which we then add to find the exact expression for any $\rho_{m}(t)$.
@@
    Probabilistic approach: We expand $e^{\mathcal{L}\Delta t}$ for small $\Delta t$ in order to connect the density operator between discretized times $j$ and $j+1$, with $t=j\Delta t$ to $t+\Delta t=(j+1)\Delta t$. Two complementary events emerge, one quantified by $d_{j}$, which is the probability to perform a jump within the time interval $\Delta t$ and $s_{j}=1-d_{j}$, quantifying the probability to remain in the respective state. Summing up all possible paths leading from the initially inverted state $\rho_{0}$ to a target state $m$ leads to an analytical expression for $\rho_{m}(t)$.
@@
    Non-Hermitian Hamiltonian approach: We start again with the recursive equations and cast it into a form involving a non-Hermitian Hamiltonian $H$. With this, the general solution involves a matrix exponential requiring a Jordan decomposition. Due to the symmetry $h_{m}=h_{\bar{m}}$, $H$ exhibits non-Hermitian degeneracies called exceptional points, which result in a modified decay compared to a non-degenerate system.
@@ II Quantum trajectories approach
$\rho(t)=\frac{1}{n_{\text{traj}}}\sum_{j=1}^{n_{\text{traj}}}\ket{\psi(t)}_{j}\bra{\psi(t)}_{j},$ (14)
@@
For a given trajectory, the system spends a total time $\tau_{m}$ in state |$\ket{m}$, with $t$ enclosed between $t_{m+1}$ (the time at which the jump from state $m+1$ into the state $m$ occurs) and $t_{m}$ (when the next jump out of the state occurs). An average of all possible trajectories then gives the value of $\rho_{m}(t)$ as the probability that the system finds itself in state $m$ at time $t$. This corresponds to an average of the segment length $\left[\theta(t-t_{m+1})-\theta(t-t_{m})\right]$ (where $\theta$ denotes the Heaviside function) over all possible values of $p_{N}$,…,$p_{1}$. The evolution is visually represented in Fig. 2. The population of the state $\rho_{m}(t)$ can then be computed as
$\rho_{m}(t)=\int_{0}^{1}\text{d}p_{N}\cdots\int_{0}^{1}\text{d}p_{m}\left[\theta(t-t_{m+1})-\theta(t-t_{m})\right],$ (15)
@@
$\theta(x)=\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\frac{\text{e}^{\text{i}\Gamma sx}}{s-\text{i}\epsilon}.$ (17)
We substitute this into Eq. (16) with the replacement $x$ by $t-\sum_{k=m+1}^{N}\tau_{k}$ first, to evaluate $\mathcal{I}_{m+1}$ and then with the replacement $x$ by $t-\sum_{k=m}^{N}\tau_{k}$ to evaluate $\mathcal{I}_{m}$. We distribute the terms with $\tau_{k}$ from the exponent into all integrals over $\tau_{k}$. This allows to compute any of the integrals which lead to the generation of poles of $s$ at $-\text{i}h_{k}$:
$\int_{0}^{\infty}\text{d}\tau_{k}e^{-\Gamma h_{k}\tau_{k}}e^{-\text{i}s\Gamma\tau_{k}}=\frac{1}{(\text{i}s+h_{k})\Gamma}.$ (18)
@@
$\mathcal{I}_{m+1}=\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\left[\prod_{k=m+1}^{N}\frac{h_{k}}{\text{i}s+h_{k}}\right]\frac{\text{e}^{\text{i}t\Gamma s}}{s-\text{i}\epsilon},$ (19)
@@
$\mathcal{I}_{m}=\frac{1}{2\pi\text{i}}\lim_{\epsilon\to 0^{+}}\int_{-\infty}^{\infty}\text{d}s\left[\prod_{k=m}^{N}\frac{h_{k}}{\text{i}s+h_{k}}\right]\frac{\text{e}^{\text{i}t\Gamma s}}{s-\text{i}\epsilon}.$ (20)
@@ III A combinatorics approach
In order to obtain the most general form of the sequences above, we first placed $r$ superoperators $B$ in the sequence (with $r$ varying from 00 to $n$) from right to left and added paddings of $C$s in the spaces between the $B$s, with $j_{N}$, $j_{N-1}$,…, $j_{N-r}$ counting the numbers of superoperators $C$ in the sequence. Summing over $r$ from 00 to $n$ and for any possible sequences (or spellings) corresponding to each fixed $r$ gives all possible combinations involved in the binomial expansion.
@@
where $\delta_{m,m^{\prime}}$ is the Kronecker delta. Summing up the infinite time series implies the need for performing the constrained sum:
@@
These kind of sums can be obtained from the expansion of given generating functions, as we describe in App. A. We make use of the result obtained there, with the identifications $n_{t}=N-m+1$ and $M=n-(N-m)$ and identifying the following generating function
@@
The constrained sum is the coefficient of the $1/z^{n-(N-m)}$ term in the series expansion of the function $f_{N-m+1}(z)$ and can be computed as the sum of the residues of $f_{N-m+1}(z)e^{z\Gamma t}$. This gives then evidently the result listed in Eq. (8).